OpenStudy (anonymous):

f(x)=x^9(x-7)^3/(x^2+6)^2 f(x)'= logarithmic differentiation

OpenStudy (anonymous):

You have y=x^9(x-7)^3/(x^2+6)^2 Now you know ln(y)=ln(x^9(x-7)^3/(x^2+6)^2) Split up the lograrithm on the right and the proceed to differentiate both sides

OpenStudy (anonymous):

ln(f(x))=ln(x^9(x-7)^3)-2ln(x^2+6) f'(x)/f(x)=9x^8(x-7)^3+3x^9(x-7)^2/x^9(x-7)^3-8x//x^2+6 Now you multiply times f(x) to each side and you already know what f(x) is

OpenStudy (anonymous):

\[In(y)=\frac{ In(x)^9(x-7)^3 }{ (x^2+6)^ 2}\]

OpenStudy (anonymous):

whats next?

OpenStudy (anonymous):

f'(x)=(9x^8(x-7)^3+3x^9(x-7)^2/x^9(x-7)^3-8x//x^2+6)x^9(x-7)^3/(x^2+6)^2

OpenStudy (anonymous):

could you put the equation by using the below bottom?

OpenStudy (anonymous):

\[\ln(f(x))=\ln(x^9(x-7)^3)-2\ln(x^2+6)\] \[f'(x)/f(x)=[(9x^8(x-7)^3+3x^9(x-7)^2)/(x^9(x-7)^3]-[4x/x^2+6)]\] \[f'(x)={[[(9x^8(x-7)^3+3x^9(x-7)^2/x^9(x-7)^3]-[4x/x^2+6)]]} x^9(x-7)^3/(x^2+6)^2\]

OpenStudy (anonymous):

Thank you, its clear! but the after x^9(x... cut off!

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

after "]]x^9"?

OpenStudy (anonymous):

yes :/

OpenStudy (anonymous):

It is the function f(x) \[x^9(x-7)^3/(x^2+6)^2\] So it is being multiplied to everything.

OpenStudy (anonymous):

is there any other easier way to do this?

OpenStudy (anonymous):

This is one of the easiest ways because you get to avoid quotient rule and it mainly involves the product rule.

OpenStudy (anonymous):

oh I see. thank you what is the final answer?

OpenStudy (anonymous):

all of it but if you notice the part where there is an\[(x-7)^2/(x-7)^3\] that could be simplified to \[1/(x-7)\] If you want to simplify further you'll probably have to multiply it out and see if anything gets factored out or cancels.

OpenStudy (anonymous):

the answer is 1/(x-7)?

OpenStudy (anonymous):

no. That was just a part that could be simplified. The answer is the long part where the x^9 was intially cut off:O

OpenStudy (anonymous):

Oh no haha x^9(x-7)^3/(x^2+6)^2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

and everything before that:O It can be a pain to write out lol

OpenStudy (anonymous):

wow, thats a really long answer!!

OpenStudy (anonymous):

would it be possible to shorten?

OpenStudy (anonymous):

It is!!! Well if you try to see if anything cancel or factors out in order to cancel

OpenStudy (anonymous):

mmm... x^9 is canceling out?

OpenStudy (anonymous):

I don't think so well maybe after a lot of expanding and factoring it will

OpenStudy (anonymous):

Ohk.

OpenStudy (anonymous):

can you help me with other question?

OpenStudy (anonymous):

Post a new question.

OpenStudy (anonymous):

Ok!